On Saturday, 63 customers at a restaurant ordered dessert and 252 customers did not order dessert. What percentage of the customers at this restaurant ordered desert on Saturday?


1
Expert's answer
2021-03-03T05:59:57-0500

"p=\frac{63}{63+252}=\frac{63}{315}=\frac{1}{5}=0.2=20\%"